OpenStudy (anonymous):

True or False. The following two lines are perpendicular. y = -x - 7 y + x = 20

OpenStudy (shinalcantara):

to know whether the lines are perpendicular or not, first thing to consider are their respective slope remember that a line perpendicular to the other is having a slope that is negative reciprocal of the other say line 1: m=1/2 line 2: m= -2

OpenStudy (anonymous):

okay

OpenStudy (shinalcantara):

slope-intercept form: y=mx + b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

okay now what do you do?

OpenStudy (shinalcantara):

for the first line, notice that it's already in the y-intercept form: y = -x - 7 that line is having slope as? m=?

OpenStudy (anonymous):

m=-x?

OpenStudy (anonymous):

or 1/2?

OpenStudy (anonymous):

-1/2?

OpenStudy (shinalcantara):

m = -1 now try to get the slope of the second line by transposing x to the right

OpenStudy (shinalcantara):

m is the value of the coefficient of x

OpenStudy (anonymous):

so x for the second one would be 1?

OpenStudy (shinalcantara):

you still need to have it in y-intercept form for you to identify the slope have it this way: y + x = 20 since the slope-intercept form is y=mx + b then subtract x from both sides of the equation y + x - x = -x + 20 y = -x + 20 what's the slope then?

OpenStudy (anonymous):

-1?

OpenStudy (shinalcantara):

yep. you have the slopes of the two lines as -1 and -1 respectively. is -1 the negative reciprocal of -1?

OpenStudy (anonymous):

yes?

OpenStudy (shinalcantara):

nope. the negative reciprocal of -1 is 1. they are parallel lines since they are having the same slope.. lines having the same slope are parallel lines. they never meet no matter how long you extend

OpenStudy (anonymous):

okay... thanks... wanna help me with another?

OpenStudy (shinalcantara):

yw :)

OpenStudy (shinalcantara):

go ahead

OpenStudy (anonymous):

True or False. The following two lines are parallel. y = 2x + 5 y = -2x